OpenStudy (anonymous):

Write the equation of a hyperbola with vertices (-5, 0) and (5, 0) and co-vertices at (0, -6) and (0, 6). Write the equation of a hyperbola with foci at (-1, 1) and (5, 1) and vertices at (0, 1) and (4, 1). Write the equation of a hyperbola with a center at (-5, -3), vertices at (-5, -5) and (-5, -1) and co-vertices at (-11, -3) and (1, -3). Write the equation of a hyperbola with foci at (3, -3) and (3, 7) and vertices at (3, -1) and (3, 5).

OpenStudy (anonymous):

please i need help.

OpenStudy (nurali):

1. Given hyperbola has a horizontal transverse axis. Its standard form: (x-h)^2/a^2-(y-k)^2/b^2=1, (h,k)=(x,y) coordinates of centre For given hyperbola: centre:(0,0) length of horizontal transverse axis=10=2a a=5 a^2=25 length of co-vertices or conjugate axis=12=2b b=6 b^2=36 Equation: x^2/25-y^2/36=1

OpenStudy (nurali):

4. (y-k)^2/a^2-(x-h)^2/b^2=1 y-coordinate of centre: mid point between y-coordinates of foci and vertices=3 x-coordinate of centre=3 centre: (3,3) length of vertical transverse axis=-1 to 5=6=2a a=3 a^2=9 foci: 2c=-3 to 7=10 c=5 c^2=25 c^2=a^2+b^2 b^2=c^2-a^2=25-9=16 b=√16=4 Equation of given hyperbola: (y-k)^2/a^2-(x-h)^2/b^2=1 (y-2)^2/9-(x-2)^2/16=1

OpenStudy (anonymous):

thnks so much... i wish i could give you my password so you can help me out